OpenStudy (friendzone):

Which is the BEST estimation of 3/400 ? A) between 7.1 and 7.2 B) between 7.2 and 7.3 C) between 7.4 and 7.5 D) between 7.3 and 7.4

OpenStudy (anonymous):

hmm one sec DAM im stupid but ill see what i get

OpenStudy (anonymous):

those numbers are multiplied by 10^-3 ?

OpenStudy (anonymous):

yeh true cool

OpenStudy (anonymous):

3/400 = (3/4) * (1/100) = 0.75 * 0.01 so it is like 7.5 but make sure first

OpenStudy (anonymous):

yeh i worked it out your correct

OpenStudy (friendzone):

Approximate 400 by finding 3/343 and 3/512 . 3/400 will be between 7 and 8. We find a more specific estimate by cubing fractional amounts between 7 and 8. 7.33 is 389.017 and 7.43 is 405.224, so 400 falls between 7.3 and 7.4.

OpenStudy (anonymous):

oh!!! 3/400 means cubic root of 400 ?! lol

OpenStudy (anonymous):

im so sorry i thought 3/400 is a fraction